Your first task is to prepare 100mL of a 1.0 mol/L solution of sodium chloride using solid sodium chloride. Calculate the mass of solid sodium chloride needed.
1
Expert's answer
2020-11-09T13:53:51-0500

m(NaCl) = M(NaCl) * Vsolution * CM = 58.5 * 0.100 * 1 = 5.85 (g)


Answer: 5.85 g of solid sodium chloride needed.